# Deterministic implementation in single-item auctions

### v1

@@ 4.1 Two buyers
An strictly increasing interim allocation profile $(\hat{x}_{1}(q),\hat{x}_{2}(q))$ is implementable by a deterministic DSIC auction if and only if
$\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))\leq q.$ (3)
@@
$\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))=q.$ (4)
The proof sketch is as follows: we first construct a corresponding three-coloring $C$ of the quantile plane, where $C\in\{0,1,2\}$, indicating that the item is allocated to no one, buyer 1 or buyer 2. Then, we rearrange a new coloring $C^{\prime}$, ensuring that the new interim allocations for both buyer 1 and buyer 2 remain unchanged. The resulting interim allocation profile can then be implemented by a deterministic DSIC auction. Furthermore, the quantile plane is partitioned by two monotone curves: $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$. Finally, we establish that $\forall q,\hat{x}_{2}(\hat{x}_{1}(q))\leq q$.
Nonequivalence of deterministic BIC and DSIC auctions with respect to interim allocations
Based on Theorem 4, we can prove the nonequivalence between deterministic BIC and DSIC auctions with respect to interim allocations. A straightforward approach is to consider the sum of interim allocations given a valuation $v$.
@@
There exists an interim allocation profile $(x_{1}(v),\cdots,x_{n}(v))$ that is implementable by a deterministic BIC auction, but there is no deterministic DSIC auction with interim allocation profile $(y_{1}(v),\cdots,y_{n}(v))$ such that for any $v$,
@@
Next, we outline the idea of the proof. We consider the two-buyer case and analyze the following randomized BIC auction:
@@
We know that the item is always sold. By Theorem 3, we have $\hat{x}_{2}(\hat{x}_{1}(q))=q$ for all $q$. This allows us to derive a series of interim allocations under a deterministic DSIC auction. However, the interim allocation may exceed 1, thereby proving the corollary.
4.2 Three buyers with constant interim allocation
@@
The key insight stems from how deterministic DSIC constraints shape the allocation space. Since deterministic DSIC requires each buyer’s allocation to depend on critical values, we can model allocations through non-overlapping regions in the valuation space. For three buyers, we partition the three-dimensional space into regions A, B, and C where buyers 1, 2, and 3 win respectively. The non-overlapping nature of these regions creates geometric constraints on their projections. Thus, we obtain a series of inequalities and derive the necessity. For sufficiency, when the condition holds, we explicitly construct deterministic allocation rules. This construction ensures DSIC while satisfying the interim allocation probabilities.
5 Discussion
In this paper, we investigated the implementation power of deterministic auctions in single-item settings, focusing on two key aspects: (revenue, welfare) pairs and interim allocation rules. Our results highlight the distinctions between deterministic BIC and deterministic DSIC auctions. Several promising directions for future research emerge from this work. First, while we identified specific distributional conditions under which deterministic DSIC auctions achieve the same (revenue, welfare) pairs as randomized auctions, it would be valuable to explore whether these conditions can be relaxed or if other classes of distributions exhibit similar properties. Second, our characterization of interim allocation rules for deterministic DSIC auctions currently applies only to two and three buyers. Generalizing this to more buyers, even under simplifying assumptions (e.g., constant interim allocations), would be a significant step forward. By addressing these open questions, future research can further illuminate the boundaries of deterministic mechanism design and its applicability across diverse economic and computational settings.
Appendix A OMITTED PROOFS FROM SECTION 3
A.1 Experimental Results Regarding Pareto Points
Only partial figure is provided here since the complete figure is too large.
A.2 Proof of Theorem 2
Proof.
$\underline{h}_{i}(q_{i})=\min_{\mathbf{q}_{-i}}h_{i}(q_{i},\mathbf{q}_{-i}).$
$\int_{0}^{1}\hat{x}_{i}^{*}(r_{i})\mathrm{d}r_{i}=\int_{s_{i}^{*}}^{1}-\underline{h}_{i}^{\prime}(r_{i})\mathrm{d}r_{i}=\underline{h}_{i}(s_{i}^{*})-\underline{h}_{i}(1)=\underline{h}_{i}(s_{i}^{*})=\int_{0}^{1}\hat{x}_{i}(r_{i})\mathrm{d}r_{i},$
$\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i})\leq\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}^{*}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i})=\int^{1}_{q_{i}=0}v_{i}(q_{i})\hat{x}_{i}^{*}(q_{i})\mathrm{d}q_{i}.$
Thus buyer $i$’s contribution to welfare weakly increases.
Lemma 5.
Proof.
$\lambda_{i,j}(q_{i})=\max\{q_{j}\mid(q_{i},\mathbf{q}_{-i})\text{ makes Border's constraint tight.}\},q_{i}\in[s_{i}^{*},1].$
  • If $\max_{i}c_{i}(q_{i})=0$, the seller keeps the item.
A.3 Proof of Lemma 1
Proof.
(a) Maximum solution
(b) Maximum solution
(c) Minimum solution
Figure 2: Optimal Solutions of Optimization Problem (2)
(a) Maximum solution
(b) Minimum solution
Figure 3: Comparison Between $P(q)$ and $\hat{P}(q)$
Put equality (6), and inequalities (7) and (8) together, we have
$\geq$~ $0.$
A.4 Proof of Lemma 3
Proof.
  • $\hat{x}(q)=0$, for $0<q\leq r_{1}$,
  • $\hat{x}(q)=k$, for $r_{1}<q\leq r_{2}$,
  • $\hat{x}(q)=q^{n-1}$, for $r_{2}<q\leq 1$.
    • If $1\notin B$, suppose that $j$ is the minimum element in $B$,
  • If $|A|=0$, $|B|>0$ and $k<\frac{r_{2}^{n-1}}{n}$:
    • If $1\in B$,
    • If $1\notin B$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
  • If $|A|=0$ and $|B|=0$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
The payment rule is according to the payment identity.
$\hat{x}_{i}(q)=Pr[q_{j}<r_{1},\forall j<i]\cdot Pr[q_{j}<r_{2},\forall j>i]\cdot Pr[q_{1}<\lambda r_{1}|q_{1}<r_{1}]=\lambda r_{1}^{i-1}r_{2}^{n-i}.$
Thus, we have
If $k<r_{2}^{n-1}$, $\hat{x}_{i}(q)=0,\forall i>1$. And we have
$\hat{x}_{1}(q)=Pr[q_{j}\leq r_{2},\forall j\neq i]\cdot Pr[q_{2}\leq\lambda r_{2}|q_{2}\leq r_{2}]=\lambda r_{2}^{n-1}\in[0,r_{2}^{n-1}].$
  • $\hat{x}(q)=0$, for $0<q\leq r_{1}$,
  • $\hat{x}(q)=q^{n-1}$, for $r_{1}<q\leq r_{2}$,
  • $\hat{x}(q)=\frac{1-r_{2}^{n}}{n(1-r_{2})}$, for $r_{2}<q\leq 1$.
  • If $|A|=0$ and $|B|=0$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
The payment rule is according to the payment identity.
For all $q<r_{1}$, $\hat{x}_{i}(q)=0$.
A.5 Proof of Lemma 2
Proof.
$\textrm{WEL}=\sum_{i=1}^{n}\int_{0}^{1}vx_{i}(v)f(v)\mathrm{d}v=\int_{0}^{1}vf(v)\sum_{i=1}^{n}x_{i}(v)\mathrm{d}v,$
$\textrm{REV}=\sum_{i=1}^{n}\int_{0}^{1}\phi(v)x_{i}(v)f(v)\mathrm{d}v=\int_{0}^{1}\phi(v)f(v)\sum_{i=1}^{n}x_{i}(v)\mathrm{d}v.$
A.6 Proof of Lemma 4
Proof.
$\mathrm{WEL}(\hat{x})=n\int_{0}^{1}v(q)\hat{x}(q)\mathrm{d}q,$
where $v(q)=F^{-1}(q)$.
Define the “reserve-type” interim allocation
$\hat{x}^{\rho}(q):=\mathbb{I}\{q>\rho\}q^{n-1}.$
Its welfare is
$W(\rho):=n\int_{\rho}^{1}v(q)q^{n-1}\mathrm{d}q,$
$0\leq k\leq k_{\max}(r_{1},r_{2}):=\frac{r_{2}^{n}-r_{1}^{n}}{n(r_{2}-r_{1})},$
and interim allocation
Let $\mathcal{M}$ be such an auction with welfare $c$. Define
$T(r):=n\int_{r}^{1}v(q)q^{n-1}\mathrm{d}q,\quad(\text{tail welfare}).$
$k(t):=\frac{c-T(r_{2}(t))}{n\int_{r_{1}}^{r_{2}(t)}v(q)\mathrm{d}q}.$
$\hat{x}_{M_{1}}(q)=\mathbb{I}\{q>\rho^{\star}\}q^{n-1}=\hat{x}^{\rho^{\star}},$
so $\mathcal{M}_{1}=\mathcal{M}^{\star}$.
A.7 The Proof of Theorem 3
Proof.
$\mathrm{WEL}(\mathcal{M}^{\max})$ $=\mathrm{WEL}(\mathcal{M}^{\min})=c,$
Appendix B OMITTED PROOFS FROM SECTION 4
B.1 Proof of Theorem 4
Proof.
Consider the plane
$A=\{(q_{1},q_{2}),0\leq q_{1}\leq 1,0\leq q_{2}\leq 1\}.$
By [nisan2007algorithmic], we know that if a DSIC auction is deterministic, then for any $i$, there exists a value $c_{i}(\mathbf{v}_{-i})$ that is independent of $v_{i}$, called the critical value, such that
Next, we rearrange the coloring $C$ to $C^{\prime}$ with fixed interim allocations.
  • $C^{\prime}(q_{1},q_{2})=1$ for all $q_{2}\leq\hat{x}_{1}(q_{1})$,
  • $C^{\prime}(q_{1},q_{2})=0$ for all $q_{2}>\hat{x}_{1}(q_{1})$ and $C(q_{1},q_{2})=1$.
  • $C^{\prime}(q_{1},q_{2})=2$ for all $q_{2}>\hat{x}_{1}(q_{1})$ and $C(q_{1},q_{2})=2$.
Figure 4: Rearrangment on three-coloring of deterministic auctions
For buyer 2, we prove that all areas that belong to buyer 2 remain the same. Otherwise, suppose that there is a point $(q_{1},q_{2})$ such that $C(q_{1},q_{2})=2$ and $C^{\prime}(q_{1},q_{2})\neq 2$. Then by the construction, it is easy to see that $q_{2}<\hat{x}_{1}(q_{1})$. By DSIC, we have $C(q_{1},q)=2$ for all $q>q_{2}$. Therefore, we have $\hat{x}_{1}(q_{1})\leq 1-\int_{q_{2}}^{1}\mathbf{I}[C(q_{1},q)=2]\mathrm{d}q_{2}=q_{2}<\hat{x}_{1}(q_{1})$, contradiction. (Here we ignore the case where $q_{2}=\hat{x}_{1}(q_{1})$ is the minimum value of $q_{2}$ such that $C(q_{1},q_{2})=2$, of which the probability is measured zero).
Do the same rearrangement for buyer 2, as shown in Figure 4, the area belongs to buyer 1 and buyer 2 are partitioned by two separate monotone curves, $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$. Since the two curves does not cross each other, (it is possible that the two curves overlap. The results still hold as the overlapped area is measured zero), we have $\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))\leq q$.
For the “if” direction, given (3), we can simply construct the DSIC auction by using the two curves, $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$, to partition the plane $A$. The condition (3) guarantees that the two curves do not cross each other, thus the construction is feasible. Moreover, if (4) holds, then the two curves fully overlap, thus all points are colored either 1 or 2. ∎
B.2 Proof of Corollary 1
Proof.
Consider the 2-buyer case and the following randomized BIC auction:
  • with probability $p$, run VCG auction.
  • with probability 1p1-p, send the item for free to buyer 1.
$\hat{x}_{1}(q)+\hat{x}_{2}(q)=2pq+1-p.$
Since
so the item is always sold. By Theorem  3, we have $\hat{x}_{2}(\hat{x}_{1}(q))=q$ for all $q$. Since $\hat{x}_{1}(0)+\hat{x}_{2}(0)=1-p$, either $\hat{x}_{1}(0)$ or $\hat{x}_{2}(0)$ must be equal to 0. Otherwise, if $\hat{x}_{1}(0)>0$ and $\hat{x}_{2}(0)>0$, then $\hat{x}_{2}(\hat{x}_{1}(0))\geq\hat{x}_{2}(0)>0$, contradiction. Thus, we without loss of generality to assume that $\hat{x}_{1}(0)=1-p$ and $x_{2}(0)=0$.
Therefore, we can get a series of sequence:
$u_{n}+u_{n+2}=2pu_{n+1}+1-p,u_{1}=0,u_{2}=1-p,$
and $\hat{x}_{1}(u_{n})=u_{n+1}$, $\hat{x}_{2}(u_{n})=u_{n-1}$.
So if the sequence $\{u_{n}\}$ does not exact reach to 1, then the interim allocation is not feasible. It is not hard to check that when $p=0.75$, $u_{5}=1.0312>1$. The corollary is proved. ∎
B.3 Proof of Theorem 5
@@
Define three valuation domain sets: $A=\{(b,c):x_{1}(*,b,c)=1\}$, $B=\{(a,c):x_{2}(a,*,c)=1\}$ and $C=\{(a,b):x_{3}(a,b,*)=1\}$. Each set represents two players’ valuation profile when the mechanism allocates the item to the remaining player.
@@
Hence $|A_{1}|=|A_{2}|$, and since $A_{1}\subseteq A_{2}$, we must have $|A_{2}\backslash A_{1}|=0$. But $A_{2}\backslash A_{1}=\{(b,c):0<cr(b,c)<1\}$. Therefore, except for a zero-measure set of $(b,c)$, we have $cr(b,c)\in\{0,1\}$, which implies $x_{1}(\cdot,b,c)$ is constant in $v$ (up to the boundary point $v=0$ or $v=1$, which does not affect interim allocations). This completes the proof. ∎

### v2

@@ 4.1 Two Buyers
A strictly increasing interim allocation profile $(\hat{x}_{1}(q_{1}),\hat{x}_{2}(q_{2}))$ is implementable by a deterministic DSIC auction if and only if
$\forall q_{1}\in[0,1],~\hat{x}_{2}(\hat{x}_{1}(q_{1}))\leq q_{1}.$ (3)
@@
$\forall q_{1},~\hat{x}_{2}(\hat{x}_{1}(q_{1}))=q_{1}.$ (4)
The proof sketch is as follows: we work on the quantile space $0,1]^{2}$ and represent any deterministic allocation rule as a three-coloring $C\in\{0,1,2\}$, where 0, 1, and 2 indicate no sale, buyer 1 wins, and buyer 2 wins, respectively. We then perform a measure-preserving rearrangement to obtain a new coloring $C^{\prime}$ that keeps each buyer’s interim allocation unchanged. Concretely, for each $q_{1}$, the measure of points colored 1 in column $q_{1}$ remains $\hat{x}_{1}(q_{1})$, and for each $q_{2}$, the measure of points colored 2 in row $q_{2}$ remains $\hat{x}_{2}(q_{2})$. This rearrangement produces a structured geometry where the plane is separated by two monotone boundary curves $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$, and each buyer’s winning set has a threshold form. In this form, monotonicity in each buyer’s own quantile is immediate, so the induced allocation can be implemented by a deterministic DSIC auction via the standard critical-threshold characterization. Finally, single-item feasibility requires the two winning regions to be disjoint, which is equivalent to the two monotone boundaries not crossing, and this gives $\forall q$, $\hat{x}_{2}(\hat{x}_{1}(q))\leq q$. The full proof of Theorem 4 is deferred to Appendix B.1.
Nonequivalence of Randomized BIC and Deterministic DSIC Auctions with Respect to Interim Allocations
Based on Theorem 4, we can prove the nonequivalence between randomized BIC and deterministic DSIC auctions with respect to interim allocations. A straightforward approach is to consider the sum of interim allocations given a valuation $v$.
@@
There exists an interim allocation profile $(x_{1}(v),\cdots,x_{n}(v))$ that is implementable by a randomized BIC auction, but there is no deterministic DSIC auction with interim allocation profile $(y_{1}(v),\cdots,y_{n}(v))$ such that for any $v$,
@@
We outline the idea of the proof. We consider the two-buyer case and analyze the following randomized BIC auction:
@@
We know that the item is always sold. By Theorem  4, we have $\hat{x}_{2}(\hat{x}_{1}(q_{1}))=q_{1}$ for all $q_{1}$. This allows us to derive a series of interim allocations under a deterministic DSIC auction. However, the interim allocation may exceed 1, thereby proving the corollary.
4.2 Three Buyers with Constant Interim Allocations
@@
Define three valuation domain sets: $A=\{(b,c):x_{1}(v_{1},b,c)=1\}$ for $\forall v_{1}$, $B=\{(a,c):x_{2}(a,v_{2},c)=1\}$ for $\forall v_{2}$ and $C=\{(a,b):x_{3}(a,b,v_{3})=1\}$ for $\forall v_{3}$. Each set represents two players’ valuation profile when the mechanism allocates the item to the remaining player.
@@
Hence $|A_{1}|=|A_{2}|$, and since $A_{1}\subseteq A_{2}$, we must have $|A_{2}\backslash A_{1}|=0$. But $A_{2}\backslash A_{1}=\{(b,c):0<cr(b,c)<1\}$. Therefore, except for a zero-measure set of $(b,c)$, we have $cr(b,c)\in\{0,1\}$, which implies $x_{1}(v,b,c)$ is constant in $v$ (up to the boundary point $v=0$ or $v=1$, which does not affect interim allocations). This completes the proof. ∎
5 Discussion
We study the implementation power of deterministic single-item auctions from two complementary perspectives: attainable (revenue, welfare) trade-offs and implementable interim allocation rules. Our results delineate a clear gap between deterministic BIC and deterministic DSIC, while also identifying distributional regimes where deterministic DSIC can match the outcomes of randomized benchmarks. Two natural directions follow. First, it remains open how far the current distributional assumptions can be weakened, and whether broader families of priors admit the same outcome equivalence. Second, our interim-allocation characterization is currently limited to two bidders (and a restricted three-bidder case). Extending it to general $n$, even under structured restrictions such as constant or low-complexity reduced forms, would substantially sharpen the boundary of deterministic implementability.
Appendix A OMITTED PROOFS FROM SECTION 3
A.1 Proof of Theorem 2
Proof.
$\underline{h}_{i}(q_{i})=\min_{\mathbf{q}_{-i}}h_{i}(q_{i},\mathbf{q}_{-i}).$
$\int_{0}^{1}\hat{x}_{i}^{*}(r_{i})\mathrm{d}r_{i}$ $=\int_{s_{i}^{*}}^{1}-\underline{h}_{i}^{\prime}(r_{i})\mathrm{d}r_{i}=\underline{h}_{i}(s_{i}^{*})-\underline{h}_{i}(1)=\underline{h}_{i}(s_{i}^{*})$
$=\int_{0}^{1}\hat{x}_{i}(r_{i})\mathrm{d}r_{i},$
$\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i})$
$\leq$ $\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}^{*}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i})$
$=$ $\int^{1}_{q_{i}=0}v_{i}(q_{i})\hat{x}_{i}^{*}(q_{i})\mathrm{d}q_{i}.$
Thus buyer $i$’s contribution to welfare weakly increases.
Lemma 5.
Proof.
$\lambda_{i,j}(q_{i})=\max\{$ $q_{j}\mid(q_{i},\mathbf{q}_{-i})\text{ makes Border's constraint}$
$\text{tight.}\},q_{i}\in[s_{i}^{*},1].$
  • If $\max_{i}c_{i}(q_{i})=0$, the seller keeps the item.
A.2 Proof of Lemma 1
Proof.
(a) Maximum solution
(b) Maximum solution
(c) Minimum solution
Figure 2: Optimal Solutions of Optimization Problem (2)
(a) Maximum solution
(b) Minimum solution
Figure 3: Comparison Between $P(q)$ and $\hat{P}(q)$
Put equality (6), and inequalities (7) and (8) together, we have
$\geq$~ $0.$
A.3 Proof of Lemma 2
Proof.
WEL $=\sum_{i=1}^{n}\int_{0}^{1}vx_{i}(v)f(v)\mathrm{d}v=\int_{0}^{1}vf(v)\sum_{i=1}^{n}x_{i}(v)\mathrm{d}v,$
REV $=\sum_{i=1}^{n}\int_{0}^{1}\phi(v)x_{i}(v)f(v)\mathrm{d}v=\int_{0}^{1}\phi(v)f(v)\sum_{i=1}^{n}x_{i}(v)\mathrm{d}v.$
A.4 Proof of Lemma 3
Proof.
  • $\hat{x}(q)=0$, for $0<q\leq r_{1}$,
  • $\hat{x}(q)=k$, for $r_{1}<q\leq r_{2}$,
  • $\hat{x}(q)=q^{n-1}$, for $r_{2}<q\leq 1$.
    • If $1\notin B$, suppose that $j$ is the minimum element in $B$,
  • If $|A|=0$, $|B|>0$ and $k<\frac{r_{2}^{n-1}}{n}$:
    • If $1\in B$,
    • If $1\notin B$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
  • If $|A|=0$ and $|B|=0$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
The payment rule is according to the payment identity.
$\hat{x}_{i}(q)=$ $Pr[q_{j}<r_{1},\forall j<i]\cdot Pr[q_{j}<r_{2},\forall j>i]$
$\cdot Pr[q_{1}<\lambda r_{1}|q_{1}<r_{1}]$
$=$ $\lambda r_{1}^{i-1}r_{2}^{n-i}.$
Thus, we have
If $k<r_{2}^{n-1}$, $\hat{x}_{i}(q)=0,\forall i>1$. And we have
$\hat{x}_{1}(q)$ $=Pr[q_{j}\leq r_{2},\forall j\neq i]\cdot Pr[q_{2}\leq\lambda r_{2}|q_{2}\leq r_{2}]$
$=\lambda r_{2}^{n-1}\in[0,r_{2}^{n-1}].$
  • $\hat{x}(q)=0$, for $0<q\leq r_{1}$,
  • $\hat{x}(q)=q^{n-1}$, for $r_{1}<q\leq r_{2}$,
  • $\hat{x}(q)=\frac{1-r_{2}^{n}}{n(1-r_{2})}$, for $r_{2}<q\leq 1$.
  • If $|A|=0$ and $|B|=0$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
The payment rule is according to the payment identity.
For all $q<r_{1}$, $\hat{x}_{i}(q)=0$.
A.5 Proof of Lemma 4
Proof.
$\mathrm{WEL}(\hat{x})=n\int_{0}^{1}v(q)\hat{x}(q)\mathrm{d}q,$
where $v(q)=F^{-1}(q)$.
Define the “reserve-type” interim allocation
$\hat{x}^{\rho}(q):=\mathbb{I}\{q>\rho\}q^{n-1}.$
Its welfare is
$W(\rho):=n\int_{\rho}^{1}v(q)q^{n-1}\mathrm{d}q,$
$0\leq k\leq k_{\max}(r_{1},r_{2}):=\frac{r_{2}^{n}-r_{1}^{n}}{n(r_{2}-r_{1})},$
and interim allocation
Let $\mathcal{M}$ be such an auction with welfare $c$. Define
$T(r):=n\int_{r}^{1}v(q)q^{n-1}\mathrm{d}q,\quad(\text{tail welfare}).$
$k(t):=\frac{c-T(r_{2}(t))}{n\int_{r_{1}}^{r_{2}(t)}v(q)\mathrm{d}q}.$
$\hat{x}_{M_{1}}(q)=\mathbb{I}\{q>\rho^{\star}\}q^{n-1}=\hat{x}^{\rho^{\star}},$
so $\mathcal{M}_{1}=\mathcal{M}^{\star}$.
A.6 The Proof of Theorem 3
Proof.
$\mathrm{WEL}(\mathcal{M}^{\max})$ $=\mathrm{WEL}(\mathcal{M}^{\min})=c,$
Appendix B OMITTED PROOFS FROM SECTION 4
B.1 Proof of Theorem 4
Proof.
Consider the plane
$A=\{(q_{1},q_{2}),0\leq q_{1}\leq 1,0\leq q_{2}\leq 1\}.$
By nisan2007algorithmic, we know that if a DSIC auction is deterministic, then for any $i$, there exists a value $c_{i}(\mathbf{v}_{-i})$ that is independent of $v_{i}$, called the critical value, such that
Next, we rearrange the coloring $C$ to $C^{\prime}$ with fixed interim allocations (see Figure 4).
  • $C^{\prime}(q_{1},q_{2})=1$ for all $q_{2}\leq\hat{x}_{1}(q_{1})$,
  • $C^{\prime}(q_{1},q_{2})=0$ for all $q_{2}>\hat{x}_{1}(q_{1})$ and $C(q_{1},q_{2})=1$.
  • $C^{\prime}(q_{1},q_{2})=2$ for all $q_{2}>\hat{x}_{1}(q_{1})$ and $C(q_{1},q_{2})=2$.
For buyer 2, we prove that all areas that belong to buyer 2 remain the same. Otherwise, suppose that there is a point $(q_{1},q_{2})$ such that $C(q_{1},q_{2})=2$ and $C^{\prime}(q_{1},q_{2})\neq 2$. Then by the construction, it is easy to see that $q_{2}<\hat{x}_{1}(q_{1})$. By DSIC, we have $C(q_{1},q)=2$ for all $q>q_{2}$. Therefore, we have $\hat{x}_{1}(q_{1})\leq 1-\int_{q_{2}}^{1}\mathbf{I}[C(q_{1},q)=2]\mathrm{d}q_{2}=q_{2}<\hat{x}_{1}(q_{1})$, contradiction. Here we ignore the case where $q_{2}=\hat{x}_{1}(q_{1})$ is the minimum value of $q_{2}$ such that $C(q_{1},q_{2})=2$, of which the probability is measured zero.
Figure 4: Rearrangment on three-coloring of deterministic auctions
Do the same rearrangement for buyer 2, as shown in Figure 4, the area belongs to buyer 1 and buyer 2 are partitioned by two separate monotone curves, $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$. Since the two curves does not cross each other, it is possible that the two curves overlap. The results still hold as the overlapped area is measured zero, we have $\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))\leq q$.
For the “if” direction, given (3), we can simply construct the DSIC auction by using the two curves, $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$, to partition the plane $A$. The condition (3) guarantees that the two curves do not cross each other, thus the construction is feasible. Moreover, if (4) holds, then the two curves fully overlap, all points are colored either 1 or 2. ∎
B.2 Proof of Corollary 1
Proof.
Consider the 2-buyer case with i.i.d. continuous priors and the following randomized BIC auction:
  • with probability $p$, run VCG auction.
  • with probability 1p1-p, send the item for free to buyer 1.
$\hat{x}_{1}(q)+\hat{x}_{2}(q)=2pq+1-p.$
Since
so the item is allocated with probability 1. By Theorem  4, we have $\hat{x}_{2}(\hat{x}_{1}(q))=q$ for all $q$. Since $\hat{x}_{1}(0)+\hat{x}_{2}(0)=1-p$, either $\hat{x}_{1}(0)$ or $\hat{x}_{2}(0)$ must be equal to 0. Otherwise, if $\hat{x}_{1}(0)>0$ and $\hat{x}_{2}(0)>0$, then $\hat{x}_{2}(\hat{x}_{1}(0))\geq\hat{x}_{2}(0)>0$, contradiction. Thus, without loss of generality, assume that $\hat{x}_{1}(0)=1-p$ and $\hat{x}_{2}(0)=0$.
Therefore, we can get a sequence:
$u_{n}+u_{n+2}=2pu_{n+1}+1-p,u_{1}=0,u_{2}=1-p,$
and $\hat{x}_{1}(u_{n})=u_{n+1}$, $\hat{x}_{2}(u_{n})=u_{n-1}$.
So if the sequence $\{u_{n}\}$ never reaches 1 exactly, then the interim allocation is not feasible. It is not hard to check that when $p=0.75$, $u_{5}=1.0312>1$. The corollary is proved. ∎